Grade 11General Physics

A car accelerates from 36 km/h to 90km/hr in 5s on straight road .what was its acceleration and how far did it travel in this time

Dhananjay
9 Years agoGrade 11

2 Answers

Kumar Karan
9 Years ago
u=36km/h=36*5/18=10m/sec
u=10m/sec
v=90km/h=90*5/18=25m/sec
v=25m/sec
t=5sec
a=(change in velocity)/t
a=25-5/5
a=4m/sec2
d=ut+1/2at2
d=10*5 + 1/2*5*5*5
d=50+50
 =100m
Mahender choudhary
7 Years ago
u=36km/hr convert into m/s.
u=36000/3600=10m/s
u=10m/s.
v=90km/hr convert into m/s.
v=90000/3600=25m/s
v=25m/s.
v=u+at
a=25-10/5
a=3m/s^2
s=ut+1/2at^2
s=10*(5)+1/2*3*(5)^2
s=87.5m